OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today? Answer 510 660 520 330

OpenStudy (anonymous):

o + t = 840 3o - 2t = 870 => 5o = 1680 + 870 => o = 2550/5 => o = 510